What is the value of the expression { 6(x - y) - 8x when x = -2 and y = 5 }
Please Explain How To Solve.

Found 2 solutions by nyc_function, tutorcecilia:
Answer by nyc_function(2741)

Replace x with -2 and y with 5 and then simplify.



6(-2 - 5) - 8(-2) =

6(-7) + 16 =

-42 + 16 = -26

Done!

Answer by tutorcecilia(2152)

{ 6(x - y) - 8x when x = -2 and y = 5 }
substitute the values for the variables:
6(-2-5)-8(-2)
solve using the Order of Operations:
6(-7)+16
-42+16
=-26